ˇ j ǰ šǧ č ŭ ǧ č ŭ č

Č ∈

L^{2} č (α· x− βt+

ϕ)

Ψ = Ae^{ı \frac{xp−t}{h̄}} (1)

∂_{x}Ψ = ıp_{x}Ψ∂_{x}^{2}Ψ = −ı/h̄^{2}· p_{x}^{2}Ψ∂_{t}Ψ = −ı/h̄· Ψ

−\frac{h̄^{2}}{2m}xΨ+Ψ = −\frac{h̄}{ı}tΨ

(2) ǧ Ψ(x, t) = ψ(x)f(t) o o memvaloroj ö memvektoroj ŭ memfunkcioj ǧ ǧ

ŭ n o_{n} ǧ

H H = ⨁_{a∈Sp(A)} H_{a}

ŭ č ekvacio laŭ la aŭ tovaloroj

O|ψ >= o|ψ >

v {e_{1}, e_{2}, …}

{c_{1}, c_{2}, …} v ǧ a_{i} v · v = v

(ov) = vo

ϕ, ψ ∈ Sϕ+ ψ ∈

S ϕ+ 0 = ϕ

∀ϕ∃−

ϕ : ϕ+

(−ϕ) =

0 cϕ =

cϕ P () nP^{n}() (Aψ)_{j} = ∑_{i} a_{ij}ψ_{i} ||Uψ|| = ||ψ|| = UψUψ = ψUUψ

∃c ∈ |||Aψ|| =

c||ψ|| Aϕψ = ϕAψ Aϕψ = ϕAψ∧ (A) =

(A) < Qψ >

(x) = xψ(x) L^{2}()

O o ǧ šǧ

P̂f(x) = f(−x)